OpenStudy (anonymous):

how should you choose two nonnegative numbers whose sum is 1 in order to maximize the product of the square of one of them and the cube of the other?

OpenStudy (raden):

we have 2 equations according information above : x + y = 1 or y = 1 - x .... (1) x^2 y^3 (assumed it has equal K) so, K = x^2y^3 ... (2)

OpenStudy (raden):

subtitute (1) to (2) K = x^2y^3 K = x^2(1-x)^3 now, can u get the first derivative (K') of K ?

OpenStudy (raden):

hint : use the produc rule, also the chain rule

OpenStudy (raden):

if K = u v then K' = vu' + uv'

OpenStudy (raden):

@syounge673

OpenStudy (anonymous):

i don't understand